(e^x)^2=e^2x prove this identity
1
Expert's answer
2020-11-17T06:18:30-0500

From basic algebra , we know that

"a^m.a^n=a^{m+n}"

Hence , "(e^x)^2=e^x.e^x=e^{x+x}=e^{2x}"